# An Information-Theoretic Approach to Identifying Formulaic Clusters in Textual Data

### v1

@@ 4.1.2 Embedding
This scope is motivated by the desire to exhaustively cover potential feature spaces while ensuring that the extracted features remain both meaningful and sufficiently frequent within the biblical corpus, thereby avoiding excessive sparsity due to the limited length of biblical texts (e.g., Antonia et al., 2014). Given the relatively small size of the corpus compared to modern datasets, this balance is particularly crucial in detecting meaningful linguistic patterns while maintaining statistical robustness.
We opt for traditional, count-based embeddings, operating under the Binomial score function (Eqs. 11-12), to ensure interpretability, particularly given the highly structured and constrained nature of biblical Hebrew. While modern deep learning techniques such as transformer-based models have revolutionized text representation in many languages (e.g., Shmidman et al., 2022), no robust pre-trained language model for biblical Hebrew currently exists. As a result, neural embeddings would require extensive domain-specific pertaining (e.g., Huertas-Tato et al., 2023), making their applicability to this study uncertain. By employing interpretable feature representations, we ensure that clustering results remain transparent and analyzable within the context of existing linguistic and source-critical scholarship.
@@ 4.1.3 Comparison With Previous Work
There is very limited prior work on unsupervised clustering approaches for exploring hypothesized divisions of biblical texts. Most previous studies have relied on supervised classification methods, where textual divisions were predefined, and models were trained using manually curated feature sets (Radday and Shore, 1985; Dershowitz et al., 2015). These approaches often rely on cherry-picked linguistic or stylistic markers that align with existing scholarly expectations, making the claim of statistical significance somewhat circular. Since the feature selection process is influenced by prior assumptions about textual divisions, such methods do not provide independent verification of whether distinct literary strata emerge naturally from the data.
By avoiding reliance on predefined features and instead allowing clusters to emerge based on intrinsic textual properties, we aim to assess whether computational methods can independently recover traditionally hypothesized partitions. Due to the lack of prior unsupervised studies, we resort to comparing our results with previous supervised approaches, evaluating whether our method aligns with existing classifications or suggests alternative structuring. In addition, we employ $k$-means clustering, previously utilized in unsupervised clustering of biblical corpora (e.g., Yoffe et al., 2023), as a baseline method to provide a comparative reference for our information-based approach. We also performed this analysis using GMM clustering for completeness, whose results are nearly identical to those of $k$-means (similarly to what was observed on synthetic data; see Fig. 1) and can be found in Appendix D.
@@ 4.2 Results
The results of our clustering analysis for the Genesis, Exodus, and Leviticus hypothesized partitions are presented in Figs. 34, and 5, respectively. These figures compare the performance of our information-based clustering method with $k$-means in identifying the hypothesized partitions of the text. To assess clustering accuracy, we categorize MCC scores into discrete performance intervals: 50-74%, 75-84%, 85-89%, 90-95%, and 96-100%, which provide a benchmark for evaluating the reliability of different parameter settings. In the case of the genealogical lists in Genesis and the P/non-P partition in Exodus, our approach resulted in a substantially larger number of parameter combinations yielding MCC scores within the two highest MCC intervals (85-89%, 90-95%) compared to $k$-means.
In the case of the Leviticus partition, while our method yielded more parameter combinations in the highest MCC interval by roughly a factor of three, $k$-means yielded slightly more MCC scores in the highest two intervals overall. This can be attributed to the fact that not all parameter combinations produce partitions with strong auto-entropy contrasts. Unlike our information-based approach, which depends on isolating structural predictability differences, $k$-means clusters based on direct feature-space separations, allowing it to identify partitions even when distinctions arise from lexical or thematic variation rather than underlying structural consistency. Because $k$-means optimizes intra-cluster similarity without requiring an information-theoretic distinction, it remains more robust across a wider range of parameter settings, whereas information-based clustering excels only when the chosen parameters successfully capture fundamental differences in textual predictability.
@@ 4.3 Formulaic Structure and Parameter Sensitivity in the P/H Partition of Leviticus
This example underscores that the identification of formulaic structure in a text is inherently dependent on the properties being examined, including grammatical and semantic scales as well as the chosen text segmentation. Different linguistic levels and contextual framing can yield divergent classifications, highlighting the need for further analysis to fully elucidate how formulaicity manifests across varying textual dimensions.
@@
Figure 6: Distinctive $n$-grams extracted from the formulaic cluster for two parameter combinations, capturing 30% of the variance (see §2.8 in Yoffe et al. (2023)). Left Panel: Clustering of morphologically-represented Leviticus using $\ell=28$, $n=4$, and $f=\text{all}$, achieving an MCC score of 94%. Right Panel: Similar to the left panel but with $\ell=20$, $n=2$, and $f=300$, achieving an MCC score of 93%. $n$-grams discussed in §4.3 as examples for H- and P-associated features are outlined in red. The insets display the self-information distributions of both clusters, with blue and orange representing the non-formulaic and formulaic clusters, respectively.
5 Conclusion
@@
We validated our method on both Gaussian-simulated and categorical datasets to assess its effectiveness across different data structures. Benchmarking against traditional clustering techniques, including k-means and various Gaussian mixture models, demonstrated that our approach provides a more stable classification of structured and unstructured components—especially in cases where the sample-to-dimension ratio is small. The information-based constraint further improved cluster separability, ensuring that formulaic structures were identified with higher precision.
@@
These findings suggest that information-theoretic clustering offers a promising avenue for unsupervised textual analysis, providing an independent measure of structural consistency that does not rely on predefined linguistic features. The stability of our method across different corpora highlights its potential applicability beyond biblical studies, particularly in fields where latent structures within textual datasets require interpretable and unbiased identification.
In future work, we will extend our framework by incorporating additional embedding techniques, such as tf-idf representations, and neural embeddings, such as transformer-based contextualized representations. These approaches will allow us to assess the model’s performance across a broader range of textual features and evaluate its adaptability to modern NLP methodologies. By integrating embeddings that capture both statistical word importance and deep semantic relationships, we aim to refine the clustering process and further improve the detection of latent structural patterns in diverse textual corpora.
@@ A.2 Resolving Power of Self-Information and Covariance Estimation
Since $\mathbb{E}[D_{M}(x)]=d$ (the expected Mahalanobis distance for a Gaussian distribution, see Appendix B), we simplify this to:
@@ A.4 Impact on Resolving Power and Signal-to-Noise Ratio (SNR)
Resolving $\text{Resolving Power}=\frac{\Delta I}{\sqrt{\text{Var}(I(x_{1}))+\text{Var}(I(x_{2}))}}.$ (37)
@@
A.5 Empirical Validation
Figure A.1: Dependence of the resolving power of self-information on sample size and dimensionality. (a) Linear dependence of variance on dimension. (b) Linear dependence of variance on $1/n$. (c) Linear dependence of $\Delta I$ on dimension.
@@ Appendix C Distinguishing Entropy-Based and Feature Distribution-Based Clusters
While the previous categorical benchmark (§3.1) demonstrated the effectiveness of our approach in distinguishing formulaic from non-formulaic structures in sparse, high-dimensional settings, an additional challenge arises when two clusters exhibit identical entropy yet differ in their feature distributions. In real-world textual corpora, stylistic or formulaic patterns are not always characterized by large entropy contrasts but may instead be defined by shifts in the distribution of certain linguistic elements. If our method is to serve as a robust tool for identifying latent structures in textual data, it must be capable of distinguishing both cases: (1) when formulaic patterns manifest as an overall reduction in entropy, and (2) when distinct compositional layers exist despite maintaining equivalent entropy levels.
@@ C.2 Experimental Results
In Fig. C.1 we compared the efficacy of feature-distribution-based clustering and formulaic clustering under these conditions. While this analysis provides insights into when each signal tends to dominate, it remains a heuristic demonstration rather than a quantitative predictive framework. We do not establish precise conditions under which one clustering paradigm will prevail over the other; developing such predictive models is an objective for future work. Instead, we illustrate that each approach becomes dominant in different parameter regimes, reinforcing the need for a nuanced understanding of structural differentiation in clustering applications.
Figure C.1: Clustering performance (MCC) as a function of key dataset parameters. Orange represents formulaic clustering, blue represents feature-distribution-based clustering. Upper Panel: MCC vs. fraction of formulaic dimensions $d_{\text{form}}$, with fixed $p_{\text{form}}=0.5$ and varying $p_{\text{feature}}\in\{0.1,0.3,0.5\}$. Middle Panel: MCC vs. baseline feature activation probability $p_{\text{feature}}$, with fixed $d_{\text{formulaic}}=0.1$ and varying $p_{\text{form}}\in\{0.8,0.5,0.1\}$ Lower Panel: MCC vs. formulaic activation probability $p_{\text{form}}$, with fixed $d_{\text{form}}=0.3$ and varying $p_{\text{feature}}\in\{0.1,0.3,0.5\}$.
As seen in Figure C.1, the relative strength of the two structural signals affects the extent to which each clustering method is effective. Formulaic clustering exhibits higher performance when the formulaic subset of dimensions is strongly expressed, suggesting that structural repetition enhances separability. Conversely, feature-distribution-based clustering improves when global differences in feature distributions are more pronounced, indicating that the method is primarily sensitive to shifts in overall distributional structure rather than localized regularities.
@@ Appendix D GMM Clustering Results
Figure D.1: Clustering results for the hypothesized partitions of the books of Genesis, Exodus, and Leviticus (left to right, respectively), using GMM clustering, similar to Fig. 3.

### v2

@@ 4.1.1 Digital Biblical Corpus and Annotation by Experts
While this study primarily employs morphological features, due to their interpretive clarity and availability in the STEPBible corpus, the method is not restricted to such representations. In settings where morphological annotation is unavailable, surface word forms can also be used. However, surface-level representations present certain challenges when evaluating formulaicity, as variation in inflection, cliticization, or syntactic context may obscure structural repetition. Morphological normalization helps reduce this variability, revealing underlying patterns of grammatical constraint. Nonetheless, there are textual domains where surface-level lexical repetition may be more meaningful.
@@ 4.1.2 Embedding
This scope is motivated by the desire to exhaustively cover potential feature spaces while ensuring that the extracted features remain both meaningful and sufficiently frequent within the biblical corpus, thereby avoiding excessive sparsity due to the limited length of biblical texts (antonia2014language, 3, e.g.,). Given the relatively small size of the corpus compared to modern datasets, this balance is particularly crucial in detecting meaningful linguistic patterns while maintaining statistical robustness.
We opt for traditional, count-based embeddings, operating under the Binomial score function (Eqs. 11-12), to ensure interpretability, particularly given the highly structured and constrained nature of biblical Hebrew. While modern deep learning techniques such as transformer-based models have revolutionized text representation in many languages (shmidman2022introducing, 56, e.g.,), no robust pre-trained language model for biblical Hebrew currently exists. As a result, neural embeddings would require extensive domain-specific pertaining (Huertas-Tato2023Oct, 27, e.g.,), making their applicability to this study uncertain. By employing interpretable feature representations, we ensure that clustering results remain transparent and analyzable within the context of existing linguistic and source-critical scholarship.
To further clarify the effect of parameter settings, we note that varying the $n$-gram size directly affects both sparsity and entropy in the feature space. Larger $n$-grams tend to emphasize rigid phraseology and reduce overlap across segments, making them useful for identifying formulaic repetitions. Smaller $n$-grams increase coverage but may conflate stylistically distinct phrases. Likewise, broader running windows ($\ell$) smooth over local variation and can help reveal segmental regularities, while narrower windows are more sensitive to fine-grained shifts. Feature count thresholds ($f$) balance interpretability and statistical stability by filtering out rare or idiosyncratic features. These parameters jointly determine which aspects of formulaicity are foregrounded and how they align with the internal segmentation of the text.
@@ 4.1.3 Comparison With Previous Work
There is very limited prior work on unsupervised clustering approaches for exploring hypothesized divisions of biblical texts. Most previous studies have relied on supervised classification methods, where textual divisions were predefined, and models were trained using manually curated feature sets radai1985genesis (47, 13). These approaches often rely on cherry-picked linguistic or stylistic markers that align with existing scholarly expectations, making the claim of statistical significance somewhat circular. Since the feature selection process is influenced by prior assumptions about textual divisions, such methods do not provide independent verification of whether distinct literary strata emerge naturally from the data.
By avoiding reliance on predefined features and instead allowing clusters to emerge based on intrinsic textual properties, we aim to assess whether computational methods can independently recover traditionally hypothesized partitions. Due to the lack of prior unsupervised studies, we resort to comparing our results with previous supervised approaches, evaluating whether our method aligns with existing classifications or suggests alternative structuring. In addition, we employ $k$-means clustering, previously utilized in unsupervised clustering of biblical corpora (yoffe2023statistical, 68, e.g.,), as a baseline method to provide a comparative reference for our information-based approach. We also performed this analysis using GMM clustering for completeness, whose results are nearly identical to those of $k$-means (similarly to what was observed on synthetic data; see Fig. 1) and can be found in D.
@@ 4.2 Results
The results of our clustering analysis for the Genesis, Exodus, and Leviticus hypothesized partitions are presented in Figures 34, and 5, respectively. These figures compare the performance of our information-based clustering method with $k$-means in identifying the hypothesized partitions of the text. To assess clustering accuracy, we categorize MCC scores into discrete performance intervals: 50-74%, 75-84%, 85-89%, 90-95%, and 96-100%, which provide a benchmark for evaluating the reliability of different parameter settings. In the case of the genealogical lists in Genesis and the P/non-P partition in Exodus, our approach resulted in a substantially larger number of parameter combinations yielding MCC scores within the two highest MCC intervals (85-89%, 90-95%) compared to $k$-means.
While each individual parameter combination in our grid (defined by $n$, $\ell$, and $f$) may yield results of interpretive interest, the primary objective of this study is to introduce and validate the clustering framework rather than to evaluate any specific configuration in detail. Accordingly, we assess method performance by analyzing the distribution of MCC scores across the entire parameter space. If a textual partition genuinely reflects differences in structural predictability, then this signal should emerge robustly across a wide range of embedding configurations, not just under fine-tuned settings.
Conversely, when only a limited subset of parameter configurations yields strong agreement with a hypothesized division, this suggests that certain levels of segmentation and feature granularity are particularly well suited to capturing certain formulaic patterns present in the data. Rather than treating this as a sign of feature selection artifacts, we interpret such behavior as evidence that distinct formulaic patterns may emerge more clearly at specific scales of representation. Some clusters, especially those linked to genre, register, or discourse framing, may only become statistically salient when, for example, the $n$-gram size and running window align with their internal structure. In this view, variation in clustering performance is not a sign of instability. Rather, it reflects both methodological sensitivity and the heterogeneous nature of formulaicity across textual layers.
This perspective is especially relevant for corpora where no predefined labels or gold standards exist. In such cases, exploring the consistency or divergence of outputs across parameter settings can help identify stable signals and locate boundaries where stylistic coherence shifts. Disagreements between parameter configurations are not necessarily failures of the method. Instead, they offer interpretive entry points into the structure of the data and the scale at which formulaic patterns operate.
In the case of the P/H partition in Leviticus, our method yielded more parameter combinations in the highest MCC interval by roughly a factor of three, but $k$-means yielded slightly more MCC scores in the highest two intervals overall. This can be attributed to the fact that not all parameter combinations produce partitions with strong auto-entropy contrasts. Unlike our information-based approach, which depends on isolating structural predictability differences, $k$-means clusters based on direct feature-space separations, allowing it to identify partitions even when distinctions arise from lexical or thematic variation rather than underlying structural consistency. Because $k$-means optimizes intra-cluster similarity without requiring an information-theoretic distinction, it remains more robust across a wider range of parameter settings, whereas information-based clustering excels only when the chosen parameters successfully capture fundamental differences in textual predictability. In §4.3, we conduct an in-depth analysis of the results for the P/H partition in Leviticus and discuss their implications.
@@ 4.3 Formulaic Structure and Parameter Sensitivity in the P/H Partition of Leviticus
Among the three case studies, the Leviticus P/H division shows the most variation in performance across parameter settings. In Genesis and Exodus, the contrasts between genealogical and narrative passages in Genesis, and between priestly and non-priestly material in Exodus, lead to many strong results clustered in the 85–90% and 90–95% MCC ranges. This suggests that the stylistic differences in those cases are captured consistently across a wide range of feature combinations. In contrast, the P/H division in Leviticus produces more extreme outcomes. It yields more high-performing results in the 90–95% range than $k$-means-based partitions, but also many more low-performing results below 75%. This behavior stands in contrast to the $k$-means baseline, which produces fewer strong results but also fewer weak ones. These patterns suggest that the formulaic differences between P and H are more sensitive to how the text is segmented and represented. In the analysis that follows, we explore how different parameter settings lead to shifts in which passages are classified as more internally regular.
@@
These results underscore that the identification of formulaic structure is not fixed to a single representation but emerges differently depending on the linguistic and structural scales chosen. In the case discussed here, distinct sets of formulaic features become salient under different $n$-gram and segmentation settings, yielding competing clusterings where either P or H is classified as more formulaic. This divergence does not undermine the method, but rather illustrates its strength: it allows researchers to test how formulaicity is distributed under different structural assumptions, exposing layered regularities that vary in grammatical, semantic, and discourse scale. The ability to generate high-confidence classifications under some parameterizations, and not others, suggests that formulaicity itself is not an absolute property, but a perspective-sensitive feature of textual organization.
@@
Figure 6: Distinctive $n$-grams extracted from the formulaic cluster for two parameter combinations, capturing 30% of the variance (see §2.8 in yoffe2023statistical (68)). Left Panel: Clustering of morphologically-represented Leviticus using $\ell=28$, $n=4$, and $f=\text{all}$, achieving an MCC score of 94%. Right Panel: Similar to the left panel but with $\ell=20$, $n=2$, and $f=300$, achieving an MCC score of 93%. $n$-grams discussed in §4.3 as examples for H- and P-associated features are outlined in red. The insets display the self-information distributions of both clusters, with blue and orange representing the non-formulaic and formulaic clusters, respectively.
5 Discussion
@@
We validated our method on both Gaussian-simulated and categorical datasets to assess its effectiveness across different data structures. Benchmarking against traditional clustering techniques, including $k$-means and various Gaussian mixture models, demonstrated that our approach provides a more stable classification of structured and unstructured components, especially in cases where the sample-to-dimension ratio is small. The information-based constraint further improved cluster separability, ensuring that formulaic structures were identified with higher precision.
@@
These findings suggest that information-theoretic clustering offers a promising avenue for unsupervised textual analysis, providing an independent measure of structural consistency that does not rely on predefined linguistic features. The stability of our method across different corpora highlights its potential applicability beyond biblical studies, particularly in fields where latent structures within textual datasets require tractable and unbiased identification.
More broadly, our findings demonstrate that structural predictability, quantified through sample-wise self-information, captures a core dimension of textual organization that aligns with, and in some cases clarifies, traditional source-critical hypotheses. In the case of computerized source criticism of the Hebrew Bible, prior computational studies have shown that blocks such as P and non-P differ in feature distributions dershowitz2015computerized (13, 68, 17). Here, we show that they also differ in their degree of formulaicity. This distinction is especially significant when made explicit through statistical modeling, as it formalizes what has often been an implicit intuition in biblical exegesis: that certain textual strata are not only lexically distinct but also stylistically more rigid, repetitive, or constrained. Crucially, our results further reveal that the expression of formulaicity is dependent on the feature set used; certain structures yield different interpretations of which passages are most internally consistent. This variability underscores the interpretive value of entropy-based methods: rather than producing a single fixed partition, they allow scholars to examine how different linguistic scales foreground different dimensions of textual structure, enabling a more nuanced and data-grounded exploration of compositional layers.
This makes entropy-based clustering particularly well suited for historical texts with compositional depth, where multiple stylistic logics may operate simultaneously. In such settings, soft clustering based on self-information variation provides a principled way to detect internal regularities without relying on predefined labels or explicit boundaries. Exploring the parameter space, across different $n$-gram granularities, segmentation schemes, and feature types, becomes an interpretive tool for probing how distinct textual regimes express constraint, formulaicity, or stylistic coherence.
For example, the distinction between Early and Late Biblical Hebrew is often expressed in systematic lexical and idiomatic shifts, making diachronic strata promising candidates for unsupervised clustering based on word-level features hurvitz1968chronological (28, 29). In the Deuteronomistic history, recurring thematic formulas and narrative structures may yield internally coherent segments whose predictability aligns with redactional layers peckham2019composition (41). Rabbinic legal corpora, such as the Mishnah, show compositional depth through repeated conditional constructions and formulaic attributions that vary across tractates or schools neusner1988mishnah (39, 18). In such cases, entropy-based clustering may help identify stylistic layers that are not organized by topic or subject matter, but by the underlying patterns of discourse structure and phrasing. Similarly, in oral-derived corpora like the Old Norse sagas, genealogical prologues, episodic transitions, and set-piece scenes create localized structural regularities that could be identified by entropy-guided segmentation clover1982medieval (10, 7). In all these contexts, the method can be used to test source-critical hypotheses, uncover latent boundaries, or evaluate competing assumptions about where and how regularity is embedded in the textual fabric.
In future work, we will extend our framework by incorporating additional embedding techniques, such as tf-idf representations, neural embeddings, and transformer-based contextualized representations. These approaches will allow us to assess the model’s performance across a broader range of textual features and evaluate its adaptability to modern NLP methodologies. By integrating embeddings that capture both statistical word importance and deep semantic relationships, we aim to refine the clustering process and further improve the detection of latent structural patterns in diverse textual corpora.
@@ A.2 Resolving Power of Self-Information and Covariance Estimation
Since $\mathbb{E}[D_{M}(x)]=d$ (the expected Mahalanobis distance for a Gaussian distribution, see B), we simplify this to:
@@ A.4 Impact on Resolving Power and Signal-to-Noise Ratio (SNR)
$\text{Resolving Power}=\frac{\Delta I}{\sqrt{\text{Var}(I(x_{1}))+\text{Var}(I(x_{2}))}}.$ (37)
@@
Figure 7 illustrates this result through empirical validation.
Figure 7: Dependence of the resolving power of self-information on sample size and dimensionality. (a) Linear dependence of variance on dimension. (b) Linear dependence of variance on $1/n$. (c) Linear dependence of $\Delta I$ on dimension.
@@ Appendix C Distinguishing Entropy-Based and Feature Distribution-Based Clusters
While the previous categorical benchmark (§3.1) demonstrated the effectiveness of our approach in distinguishing formulaic from non-formulaic structures in sparse, high-dimensional settings, an additional challenge arises when two clusters exhibit identical entropy yet differ in their feature distributions. In real-world textual corpora, stylistic or formulaic patterns are not always characterized by large entropy contrasts but may instead be defined by shifts in the distribution of certain linguistic elements. If our method is to serve as a robust tool for identifying latent structures in textual data, it must be capable of distinguishing between two cases: (1) when formulaic patterns manifest as an overall reduction in entropy, and (2) when distinct compositional layers exist despite maintaining equivalent entropy levels.
@@ C.2 Experimental Results
In Figure 8, we illustrate the comparison between the efficacy of feature-distribution-based clustering and formulaic clustering under these conditions. While this analysis provides insights into when each signal tends to dominate, it remains a heuristic demonstration rather than a quantitative predictive framework. We do not establish precise conditions under which one clustering paradigm will prevail over the other; developing such predictive models is an objective for future work. Instead, we illustrate that each approach becomes dominant in different parameter regimes, reinforcing the need for a nuanced understanding of structural differentiation in clustering applications.
Figure 8: Clustering performance (MCC) as a function of key dataset parameters. Orange represents formulaic clustering, blue represents feature-distribution-based clustering. Upper Panel: MCC vs. fraction of formulaic dimensions $d_{\text{form}}$, with fixed $p_{\text{form}}=0.5$ and varying $p_{\text{feature}}\in\{0.1,0.3,0.5\}$. Middle Panel: MCC vs. baseline feature activation probability $p_{\text{feature}}$, with fixed $d_{\text{formulaic}}=0.1$ and varying $p_{\text{form}}\in\{0.8,0.5,0.1\}$ Lower Panel: MCC vs. formulaic activation probability $p_{\text{form}}$, with fixed $d_{\text{form}}=0.3$ and varying $p_{\text{feature}}\in\{0.1,0.3,0.5\}$.
As seen in Figure 8, the relative strength of the two structural signals affects the extent to which each clustering method is effective. Formulaic clustering exhibits higher performance when the formulaic subset of dimensions is strongly expressed, suggesting that structural repetition enhances separability. Conversely, feature-distribution-based clustering improves when global differences in feature distributions are more pronounced, indicating that the method is primarily sensitive to shifts in overall distributional structure rather than localized regularities.
@@ Appendix D GMM Clustering Results
In Figure 9, we illustrate the classification results using the Gaussian Mixture Model (GMM) algorithm as applied to the three biblical texts, as discussed in §4.
Figure 9: Clustering results for the hypothesized partitions of the books of Genesis, Exodus, and Leviticus (left to right, respectively), using GMM clustering, similar to Fig. 3.